OpenStudy (anonymous):

write an equation in slope intercept, point slope, or standard form for the line with given information. Slope 2, y-intercept -4

OpenStudy (anonymous):

y = mx +b m = slope b = y-intercept

Nnesha (nnesha):

slope intercept form equation is y = mx + b point slope is ( y - y1) = m ( x - x1) standard form Ax + Bx = c

OpenStudy (godlovesme):

Point-slope formula is y-y1= m(x-x1) Slope-intercept y=MX+b Standard ax+b=c

OpenStudy (anonymous):

thanks. i can't believe i forgot how to do this lol

Nnesha (nnesha):

did you get it now ???

Nnesha (nnesha):

y = mx + b b is y-intercept which is -4 slope is m which is 2 so plug that into y = mx + b